# Predicting Chemically Accurate Adsorption Energy Using an Interpretable Deep Learning Model Pretrained by GGA Calculation Data

### v1

@@ Results
Building upon the robust performance and transferability of DOTA-PBE, we further fine-tuned the model to achieve chemically accurate predictions of adsorption energies for specific adsorbates across metallic and intermetallic surfaces. The notable scarcity of chemically accurate adsorption energy data requires a model architecture with a minimum requirement of labeled chemically accurate adsorption energy data to fine-tune the model. Taking CO adsorption as an example, we first incorporated 988 LDOS-CO adsorption energy pairs (Table S2) across different sites of 174 metallic and intermetallic surfaces, evaluated by PBE, RPBE [88], and PBEsol [89] functionals, respectively. Data derived from different functionals expanded the perceptive chemical space, thereby enhancing the ability of the model to learn orbital interaction patterns between CO and metallic surfaces. Importantly, the foundation of PBE-level orbital interaction understanding, combined with GGA-level insights into CO-surface interactions, enabled exceptional data efficiency to achieve chemical accuracy predictions with only four high-fidelity data points. Each high-fidelity data point comprised experimentally measured CO adsorption energies (Pd(111), Cu(111), Ir(111), or Ni(111)) through precise microcalorimetry, paired with corresponding PBE-calculated surface LDOS profiles and HSE06-calculated gaseous CO LDOS characteristics (Fig. 4a). We compared adsorption energies of CO over Pt(111) and Rh(111) (Fig. 4c) predicted by the DOTA model with different input functional combinations for surface and adsorbate LDOS profiles with standard PBE, PBEsol, HSE06, RPA, and experimental results. As introduced from Blyholder’s model, [90] CO adsorption is primarily formed through the -donation between CO HOMO (5 occupied orbital) and the metal conduction bands and the d- backbonding between CO LUMO (2 unoccupied orbital) and the metal valence bands. Accordingly, the correct CO adsorption energy and binding sites rely on the correct description of the electronic structures of CO and the metallic surfaces. As reported previously, [91] HSE06 functional is poor at predicting CO adsorption because it fails to describe the delocalized electronic structures of metallic surfaces, while the underestimated HOMO-LUMO gap of CO by PBE and PBEsol functionals [92, 93] should account for the overbinding of CO over Pt(111) and Rh(111) surfaces. [94, 95] Moreover, these deficiencies result in the mistakenly predicted fcc sites for CO chemisorption over these two surfaces. This manifests that the underestimated HOMO-LUMO gap by PBE functional is mainly responsible for the "CO puzzle", which was speculated by previous reports. Interestingly, despite the accurate CO adsorption energy, neither the CO HOMO-LUMO gap nor the metal cohesive energies could be accurately calculated by RPA (Fig. 4), which is consistent with previously reported results. [58] Importantly, HSE06 exhibits the best predictive performance for CO HOMO-LUMO energy levels, while PBE functional excels at the prediction of metal cohesive energy, indicating that HSE06 and PBE functionals are good at the description of the molecular orbitals and metal surfaces, respectively. Consequently, the input of HSE06-calculated gaseous CO LDOS characteristics and PBE-calculated surface LDOS profiles (HSE06/PBE) could provide superior results, with the predicted adsorption energy differing by less than 0.1 eV from the experimental values. More importantly, the model with the HSE06/PBE input combination could correctly predict the top site as the preferential binding site for CO adsorption, aligning with the experimental results. [18] The accuracy of the multimodal integration capitalizes on the advantages of the experimental precision of microcalorimetry in adsorption energy, the reliability of PBE functional for the description of electronic structures of metallic surfaces [96], and the superior performance of HSE06 beyond RPA for CO molecular orbital description [97, 84].
@@ Discussion
The DOTA model integrates LDOSs from both the surface and the molecule as input. Different from the d-band center and Fermi level theories, this approach enables accurate prediction of adsorption energies across diverse adsorbates (MAE = 0.066 eV, MAPE = 1.53) without reliance on predefined theoretical assumptions. Leveraging a Transformer architecture and integrated gradients algorithm, DOTA provides interpretable insights into the contribution of individual energy levels to adsorption strength. Established upon two key physical insights of the origin of the adsorption strength from orbital interactions and the functional-independent DOS-adsorption relationship, DOTA demonstrates its exceptional transferability. This permits integration of experimental sources and DFT data from different functionals and calculation settings. It could further achieve chemical accuracy in predicting molecular adsorption energies across extensive metallic and intermetallic surfaces, primarily trained on GGA-level datasets augmented by few-shot high-fidelity experimental adsorption energy data. Thus, DOTA represents a transformative leap forward in computational catalysis by addressing longstanding challenges such as functional dependency, limited availability of high-fidelity data, and the need for interpretable machine learning models. By synergizing deep learning architectures and physically informed features like LDOS, it achieves unprecedented accuracy and generalizability across chemical environments. Its capacity to unify multi-fidelity data sources and deliver mechanistic insights through interpretable attention paradigms establishes DOTA as a versatile framework for bridging experimental precision with computational efficiency. This could facilitate the construction of a large dataset of the adsorption energies with high fidelity, accelerating the discovery of next-generation catalysts and electrode materials.
@@ Methods
All DFT calculations were performed with Vienna ab initio Simulation Package (VASP) [98, 99] with periodic boundary and the projector augmented wave (PAW) pseudopotentials [98, 100]. The Catalysis Kit (CatKit) software [44] and Atomic Simulation Environment (ASE) Python library [101, 102] were used for generating and classifying the adsorption structures. The chemisorption structures, except for CO adsorption, were obtained in reference [45] for DOTA model pretraining, where PBE functionals were used for both adsorption energy and DOS profile calculations. For CO adsorption, RPBE functional and PBEsol functional were also used for adsorption energy and DOS profile calculations, while the hybrid HSE06 functional was used only for electronic structure calculations. All calculations were conducted with spin polarization, using a kinetic energy cutoff of 450 eV. The p(22) 3-layer slab was employed for each metallic and intermetallic surface. The corresponding Monkhorst-Pack k-point grid [103] was 771, with a first-order Methfessel-Paxton smearing of 0.1 eV. The resolution of calculated lm- and site-DOSs with reference to the Fermi level, which were employed for DOTA training, was 0.01 eV.
The DOTA model was implemented with the PyTorch Python library [104]. The AdamW optimizer with an initial learning rate of 0.0002 and the L1 loss were used for training. An exponential cosine annealing scheduler was implemented to adjust the learning rate, and 150 epochs were used, with 50 epochs for warming up, and a batch size of 256. The contribution of each energy level to the adsorption energy was calculated using the interpretable integrated gradients algorithm implemented in the Captum library [87] with default parameters.
@@
Code for DOTA is available as an open source repository on GitHub, https://github.com/zhzhang321/DOTA
@@
The authors declare no competing interests.
Acknowledgements
This work was supported by the National Key Research and Development Program of China (2023YFA1507601 and 2021YFA1500700). We thank the Center for High Performance Computing at Shanghai Jiao Tong University for providing the computing resources of the Siyuan-1 cluster.
Author contributions
Z.Zhang wrote the code and performed theoretical simulations. X.-M.C. designed the study, analyzed the data, and wrote the paper.

### v2

@@ Results
Building upon the robust performance and transferability of DOTA-PBE, we further fine-tuned the model employing the multi-head strategy to predict chemically accurate adsorption energies for specific adsorbates across metallic and intermetallic surfaces. The notable scarcity of chemically accurate data requires a model architecture with a minimum requirement of labelled chemically accurate data to fine-tune the model. Taking CO adsorption as an example, we first incorporated 988 LDOS-(CO) pairs (Table S2) across different sites of 174 metallic and intermetallic surfaces, evaluated by PBE, RPBE [88], and PBEsol [89] functionals, respectively, as one fine-tuning dataset. Data derived from different functionals expanded the perceptive chemical space, thereby enhancing the ability of the model to learn orbital interaction patterns between CO and metallic surfaces. Importantly, the foundation of PBE-level orbital interaction understanding, combined with GGA-level insights into the interactions between CO and the surface, enabled exceptional data efficiency to achieve chemical accuracy predictions with only four high-fidelity data points in the other fine-tuning dataset. Each high-fidelity data point comprised experimentally measured CO adsorption energies ((exp.) [18]: Pd(111), Cu(111), Ir(111), or Ni(111)) as output, paired with corresponding input PBE-calculated surface LDOS profiles (LDOS(surf, PBE)) and HSE06-calculated gaseous CO LDOS (LDOS(CO(g), HSE06)) characteristics (Fig. S1a). We compared (CO) over Pt(111) and Rh(111) (Fig. 4a and Fig. 4b) predicted by the DOTA model with different input functional combinations for adsorbate/surface LDOS profiles, the standard PBE, PBEsol, HSE06, RPA, and experimental results. As introduced from Blyholder’s model [90], CO adsorption is primarily formed through the -donation between CO HOMO (5 occupied orbitals) and the metal conduction bands and the d- backbonding between CO LUMO (2 unoccupied orbitals) and the metal valence bands. Accordingly, the correct (CO) and binding sites rely on the correct description of the electronic structures of CO and the metallic surfaces. Fig. 4a shows that both the standard PBE and PBE/PBE combinations results in CO overbonding and mistakenly predicted fcc sites for CO chemisorption over these two surfaces. More poor results were predicted by the standard PBESol and PBESol/PBESol combinations. It was consistent with the previous report [91, 92] that the significantly underestimated HOMO-LUMO gap of CO by PBE and PBEsol functionals [93, 94] should account for the overbinding of CO over the Pt(111) and Rh(111) surfaces. Nevertheless, PBE exhibit good performance at the description of the metallic surface electronic structure, as demonstrated by the closer metal cohesive energy to the experimental value (Fig. 4a and Fig. 4b). The standard HSE06 functional is also poor at predicting CO adsorption because it fails to describe the delocalised electronic structures of metallic surfaces. [95, 84] Interestingly, despite the accurate (CO), neither the CO HOMO-LUMO gap nor the metal cohesive energies could be accurately calculated by RPA (Fig. S1b, Fig. 4a, and Fig. 4b), which is consistent with previously reported results. [58] Importantly, HSE06 exhibits the best predictive performance for CO HOMO-LUMO energy levels, while the PBE functional excels at the prediction of metal cohesive energy [58], indicating that the HSE06 and PBE functionals are good at the description of the molecular orbitals and metal surfaces, respectively. Consequently, the input of LDOS(CO(g), HSE06) and LDOS(surf, PBE) (HSE06/PBE) could provide superior results, with the predicted differing by less than 0.04 eV from the experimental values. More importantly, the model with the HSE06/PBE input combination could correctly predict the top site as the preferential binding site for CO adsorption, aligning with the experimental results. [18] This also manifests that the underestimated HOMO-LUMO gap by the PBE functional is mainly responsible for the ”CO puzzle”, which was speculated by previous reports. [96]. Moreover, the accuracy of the multimodal integration capitalises on the advantages of the experimental precision of microcalorimetry in , the reliability of the PBE functional for the description of electronic structures of metallic surfaces [97], and the superior performance of HSE06 beyond RPA for CO molecular orbital description [95, 84].
Encouraged by the chemically accurate prediction for CO, we further examined the transferability of the fine-tuned DOTA framework to dissociative adsorption of and , under extremely data-limited regimes. For these molecules, the was computed as , where (X = H or O) was predicted by DOTA from atomic LDOS inputs, and or D(O=O) denotes the standard gas-phase bond dissociation energy [98]. Due to the scarcity of high-fidelity experimental dissociative adsorption energies, the model was trained on only 3 data points for H and 2 for O, and validated against 5 independent measurements for each. As shown in Fig. 4, predictions for both training and validation sets cluster closely around the parity line across all three molecules based on minimal supervision. Moreover, the MAE value of 0.033 eV is within the chemical accuracy. These preliminary results suggest that DOTA’s orbital-aware architecture enables robust few-shot learning not only for molecular adsorption but also for dissociative adsorption, highlighting its potential for modelling more surface reactions such as hydrogen evolution or oxygen reduction.
@@ Discussion
Established upon two key physical insights, the orbital-aware adsorption strength and the functional-independent LDOS-adsorption relationship, the DOTA model permits the alignment of experimental sources and multi-fidelity DFT data. It could further achieve chemical accuracy in predicting molecular adsorption energies across extensive metallic and intermetallic surfaces, primarily trained on GGA-level datasets augmented by minimum high-fidelity experimental data. This demonstrates its exceptional transferability among elements. Thus, DOTA represents a transformative leap in computational catalysis by addressing longstanding challenges such as functional dependency, limited availability of high-fidelity data, and the need for interpretable machine learning models. By synergising deep learning architectures and physically informed features like LDOS, it achieves unprecedented accuracy and generalisability across chemical element environments.
Herein, the limited high-fidelity experimental data from the (111) surface render the DOTA model to be trained upon the data from the flat (111) surface. Because the orbital interaction patterns between the adsorbate and metallic surface would vary with the exposed facet, the DOTA model would be sensitive to the surface structures. In addition, the available high-fidelity experimental data correspond to a specific surface coverage, necessitating consistency with the GGA data for fine-tuning. Nevertheless, with the rapid accumulation of GGA data, the capacity of DOTA to unify multi-fidelity data sources enables the rapid construction of a large dataset of adsorption energies with high fidelity over different exposed surfaces, requiring an extremely small quantity of extra high-fidelity experimental or high-ladder quantum chemistry data. This could significantly accelerate the discovery of next-generation catalysts and electrode materials.
@@ Methods
All the DFT calculations were performed with Vienna ab initio Simulation Package (VASP) [99, 100] with periodic boundary and the projector augmented wave (PAW) pseudopotentials [99, 101]. The Catalysis Kit (CatKit) software [44] and the Atomic Simulation Environment (ASE) Python library [102, 103] were used to generate and classify the adsorption structures. The chemisorption structures were obtained in reference [45] for DOTA model pretraining, where the PBE functional was used for both and DOS profile calculations. For the fine-tuning stage, PBE, RPBE and PBEsol functionals were used for and DOS profile calculations. The hybrid HSE06 functional was used only for the DOS profile of the gaseous adsorbate. All calculations were conducted with spin polarisation, using a kinetic energy cutoff of 450 eV. The p(22) 3-layer slab was employed for each metallic and intermetallic surface. The corresponding Monkhorst-Pack k-point grid [104] was 771, with a first-order Methfessel-Paxton smearing of 0.1 eV. The resolution of calculated lm- and site-DOSs with reference to the Fermi level, which were employed for DOTA training, was 0.01 eV.
The DOTA model was implemented with the PyTorch Python library [105]. The AdamW optimiser with an initial learning rate of 0.0002 and the L1 loss were used for training. An exponential cosine annealing scheduler was implemented to adjust the learning rate, with 150 epochs used, comprising 50 epochs for warming up. A batch size of 256 was employed. The contribution of each energy level to the was calculated using the interpretable integrated gradients algorithm implemented in the Captum library [87] with default parameters.
@@
Code for DOTA is available as an open-source repository on GitHub, https://github.com/zhzhang321/DOTA
Acknowledgements
This work was supported by the National Key Research and Development Program of China (2023YFA1507601 and 2021YFA1500700). We thank the Centre for High Performance Computing at Shanghai Jiao Tong University for providing the computing resources of the Siyuan-1 cluster.
Author information
Author contributions
Z.Zhang wrote the code, performed theoretical simulations, and wrote the paper. X.-M.C. designed the study, analysed the data, and wrote the paper. Both authors discussed the results and improved the manuscript.
Corresponding authors
Correspondence to Xiao-Ming Cao.
Ethics declarations
@@ Acknowledgements
All authors declare no competing interests.
Supplementary information
Supplementary Figs. 1-3 and Tables 1-4.
Rights and permissions
Reprints and permissions information is available at http://www.nature.com/reprints.